OpenStudy (samsan9):

The vertical asymptote of the function y=log(x-2)+10 is:? A.) x=2 B.) x=10 C.)x=-2 D.) x=-10

OpenStudy (anonymous):

Try plugging in the given \(x\) values and see what happens.

OpenStudy (samsan9):

2would make the inside 0 and 10 would make 8 -2 would make it -4 and -10 would make it -12

OpenStudy (anonymous):

Right, but consider the domain of the logarithm. For any base \(b\), \(\log_b f(x)\) is only defined when \(f(x)>0\). This means that whatever the "inside" of the logarithm is, it must be greater than zero. This also means asymptotes occurs for whichever values of \(x\) make the inside function 0. Which value of \(x\) would it be for \(\log(x-2)\)?

OpenStudy (samsan9):

-2 since it would make it less than 0?

OpenStudy (anonymous):

-2 would make it less than 0, but an asymptote occurs for the \(x\) that makes it exactly 0. For example, \(\ln x\) is undefined at \(x=0\) because the inside function, \(x\) is 0 when \(x=0\).

OpenStudy (samsan9):

oh wait my bad it would 10

OpenStudy (anonymous):

Not 10 either, since \[\log(10-2)+10=\log8-2\] is defined. What makes \(x-2\) equal to 0?

OpenStudy (samsan9):

so just 2 ?

OpenStudy (anonymous):

Yes, when \(x=2\) you have \(\log(2-2)+10=\log0+10\), but \(\log0\) is undefined.